A book, a hammer, and a lamp are weighed and the results are recorded. The book and the hammer together weigh 8 pounds. The hammer and the lamp together weigh 10 punds. The book and the lmap together weigh 6 punds. How heavy, in pounds is the lamp?
Answer by scott8148(6628)

b + h = 8

h + L = 10

b + L = 6

adding 2nd and 3rd equations ___ h + b + 2L = 16

subtracting 1st equation ___ 2L = 8